# Deep End-to-end Adaptive k-Space Sampling, Reconstruction, and Registration for Dynamic MRI

### v1

@@ 0.C.2 Addiational Figures
(a) Case I: Original setup, joint loss ($\alpha=1,\beta=1$).
@@
(b) Case I: Decoupled loss setup ($\alpha=1,\beta=1$).
@@
(c) Case I: Joint loss ($\alpha=1,\beta=3$).
@@
(d) Case II: Original setup, joint loss ($\alpha=1,\beta=1$).
@@
(e) Case II: Decoupled loss setup ($\alpha=1,\beta=1$).
@@
(f) Case II: Original setup, joint loss ($\alpha=1,\beta=3$).
Figure 13: Example results for two cases, shown at various temporal frames ($\tau$) relative to the reference image, for different loss function setups at $R=8$.
@@
Figure 14: Example results for three cases, for shown at various temporal frames ($\tau$) relative to the reference image, for different choices of reconstruction network at different accelerations.
Refer to captionRefer to caption
(a) Original setup.
Refer to captionRefer to caption
(b) Original setup with larger model.
Refer to captionRefer to caption
(c) Voxelmorph.
Refer to captionRefer to caption
(d) Transmorph.
Refer to captionRefer to caption
(e) Optical Flow ILK.
Refer to captionRefer to caption
(f) DEMONS.
Figure 15: Example results I for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of registration network at $R=8$.
Refer to captionRefer to caption
(a) Original setup.
Refer to captionRefer to caption
(b) Original setup with larger model.
Refer to captionRefer to caption
(c) Voxelmorph.
Refer to captionRefer to caption
(d) Transmorph.
Refer to captionRefer to caption
(e) Oprical Flow ILK.
Refer to captionRefer to caption
(f) Optical Flow TVL1.
Figure 16: Example results II for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of registration network at $R=4$.
Refer to captionRefer to caption
(a) Original setup.
Refer to captionRefer to caption
(b) Original setup with larger model.
Refer to captionRefer to caption
(c) Voxelmorph.
Refer to captionRefer to caption
(d) Transmorph.
Figure 17: Example results III for an aorta case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of registration network at $R=6$.
@@
Figure 18: Example results I for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of adaptive sampling at $R=8$.
@@
Figure 19: Example results II for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of adaptive sampling at $R=6$.

### v2

@@ C.3 Additional Figures
(a) Original setup.
@@
(b) Voxelmorph.
@@
(c) Transmorph.
@@
(d) Optical Flow ILK.
@@
(e) DEMONS.
Figure C1: Example results I for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of registration network at $R=8$.
@@
(a) Original setup.
Refer to captionRefer to caption
(b) Voxelmorph.
Refer to captionRefer to caption
(c) Transmorph.
Refer to captionRefer to caption
(d) Oprical Flow ILK.
Refer to captionRefer to caption
(e) Optical Flow TVL1.
Figure C2: Example results II for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of registration network at $R=4$.
Refer to captionRefer to caption
(a) Original setup.
Refer to captionRefer to caption
(b) Voxelmorph.
Refer to captionRefer to caption
(c) Transmorph.
Figure C3: Example results III for an aorta case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of registration network at $R=6$.
@@
Figure C4: Example results for three cases, for shown at various temporal frames ($\tau$) relative to the reference image, for different choices of reconstruction network at different accelerations.
@@
Figure C5: Example results I for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of adaptive sampling at $R=8$.
@@
Figure C6: Example results II for a cardiac cine case, shown at various temporal frames ($\tau$) relative to the reference image, for different choices of adaptive sampling at $R=6$.
Refer to captionRefer to caption
(a) $\alpha=1$, $\beta=1$ (original setup)
Refer to captionRefer to caption
(b) $\alpha=1$, $\beta=2$
Refer to captionRefer to caption
(c) $\alpha=1$, $\beta=3$
Refer to captionRefer to caption
(d) $\alpha=\frac{3}{2}$, $\beta=2$
Refer to captionRefer to caption
(e) $\alpha=0$, $\beta=1$
Refer to captionRefer to caption
(f) $\alpha=2$, $\beta=1$
Figure C7: Example I results for a case, shown at various temporal frames ($\tau$) relative to the reference image, for different joint loss weighting parameters at $R=8$.
Refer to captionRefer to caption
(a) $\alpha=1$, $\beta=1$ (original setup)
Refer to captionRefer to caption
(b) $\alpha=1$, $\beta=2$
Refer to captionRefer to caption
(c) $\alpha=1$, $\beta=3$
Refer to captionRefer to caption
(d) $\alpha=\frac{3}{2}$, $\beta=2$
Refer to captionRefer to caption
(e) $\alpha=0$, $\beta=1$
Refer to captionRefer to caption
(f) $\alpha=2$, $\beta=1$
Figure C8: Example II results for a case, shown at various temporal frames ($\tau$) relative to the reference image, for different joint loss weighting parameters at $R=8$.
Refer to captionRefer to caption
(a) Joint training.
Refer to captionRefer to caption
(b) Decoupled training
Refer to captionRefer to caption
(c) Joint training.
Refer to captionRefer to caption
(d) Decoupled training
Figure C9: Example results for two cases, shown at various temporal frames ($\tau$) relative to the reference image, comparing joint vs decoupled training setups at $R=8$.
Refer to captionRefer to caption
(a) All loses.
Refer to captionRefer to caption
(b) $\mathcal{L}_{\text{reg}}=\mathcal{L}_{\text{sim}}=\mathcal{L}_{\text{ssim2D}}$
Refer to captionRefer to caption
(c) $\mathcal{L}_{\text{reg}}=\mathcal{L}_{\text{sim}}=\mathcal{L}_{1}$
Refer to captionRefer to caption
(d) No smooth loss.
Figure C10: Example results for a case, shown at various temporal frames ($\tau$) relative to the reference image, comparing different loss choice setups at $R=8$.